# Blind and Topological Interference Managements for Bistatic Integrated Sensing and Communication

### v4

@@ VI Simulation Results
Figure 9: Simulation results: (a) BIA (b) TIM.
We simulate proposed schemes for the bistatic ISAC systems with heterogeneous coherent times and the interference channel in Example 2, and with heterogeneous connectivity and interference channel in Example 3, compared with the method by treating interference as noise (TIN) at the sensor. [^9] Since the sensor channel is unknown to the sensor, the successive interference cancellation (SIC) decoding strategy in [38] which first decodes the communication messages and then estimates the sensing state, cannot be used in our simulation.
@@ VI-A Heterogeneous Coherence Times
[…]
@@ VI-B Heterogeneous Connectivity
We modify the simulation in the above case to the interference channel with partial connection as in Example 3. Note that in this case, the transmitter only knows the network topology, without known the CSI of any receiver nor sensor. Our method provides a gain of [3.87, 18.892] dB on the estimation accuracy as shown in Fig. 9(b).
@@ VII Conclusion
This paper explored the interference management issue in bistatic ISAC systems, specifically how to effectively cancel the interference from the communication signals to the sensor, by applying the BIA and TIM strategies. The key is to leverage the heterogeneity (on the coherence times or on the connectivity) between the sensor and receivers. By using the communication and sensing degrees of freedom as the metric, the proposed schemes provide improved tradeoff points, compared to the time-sharing between the communication-only and sensing-only points. Simulations on the proposed schemes and the existing TIN scheme are provided, showing that the proposed schemes reduce the sensing CEE, while maintaining the same communication performance.
@@ Appendix A Proof of Theorem 3: MU-MIMO
Appendix B Proof of Theorem 4-b: Topology 2 with Non-cooperative Transmitters
This is the general description of the proposed TIM scheme for Theorem 4-b. For the Topology 2 based on the $({\mathsf{K}},U,D)$ neighboring antidotes network, we propose an ISAC scheme based on TIM achieving the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{1}{{\mathsf{K}}-D+U},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}\right)$ in (5). Each receiver is disconnected from the previous $U$ and next $D\geq U$ transmitters, and connected to other ${\mathsf{K}}-U-D>0$ transmitters. Remind that the sensor is a subset of one communication that the network topology could be pressed as
@@ Appendix B Proof of Theorem 4-b: Topology 2 with Non-cooperative Transmitters
for each $i,k\in\{1,\ldots,{\mathsf{K}}\}$, To simplify notation without loss of generality, we set $i=1$, thus we can get ${\mathcal{R}}_{s}=\{2+D\,\ldots,{\mathsf{K}}-U\}$. Let $t_{0}=({\mathsf{K}}-D-U)({\mathsf{K}}-D+U)$, during these $({\mathsf{K}}-D-U)({\mathsf{K}}-D+U)$ time slots, the sensor will obtain ${\mathsf{K}}-D-U$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{{\mathsf{K}}-D-U}^{[0]}\in{\mathbb{C}}^{({\mathsf{K}}-D-U)\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{{\mathsf{K}}-D-U}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{{\mathsf{K}}-D-U}^{[0]}]^{H}=I$. In each period of $({\mathsf{K}}-D+U)$ time slots, we let each receiver decode $U+1$ communication messages, and let the sensor obtain one effective observation. In the next, we will illustrate our proposed scheme for the first $K-D+U$ time slots, where each receiver $k\in\{1,2,...,{\mathsf{K}}\}$ should recover $U+1$ messages $W_{1}^{[k]},\ldots,W_{U+1}^{[k]}\in{\mathbb{C}}$.
@@
for each $i\in{\mathcal{R}}_{s}\}$, where ${\mathbf{x}}_{s}^{[j]},j\in\mathcal{R}_{s}$ are the sensing signals to be determined later. For the transmitters not connected to the sensor, design based on the cyclic coding,
@@
where $\max(a,b)$ is the larger value (i.e., $a$ if $a\geq b$, otherwise $b$) and $\min(a,b)$ is the smaller value (i.e., $a$ if $a\leq b$, otherwise $b$). Therefore, the dimension of the interference signals is aligned to ${\mathsf{K}}-D-1$, by the linearly independece of ${\mathsf{K}}-D-1$ vectors ${\mathbf{v}}^{[<k+D+1>_{{\mathsf{K}}+1}]}\ {\mathbf{v}}^{[<k+D+2>_{{\mathsf{K}}+1}]}\ \ldots{\mathbf{v}}^{[<k+{\mathsf{K}}>_{{\mathsf{K}}+1}]}$, receiver $k$ can recover $U+1$ messages from the first ${\mathsf{K}}-D+U$ time slots. Thus, the communication degree of freedom is $\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}$.
@@
$=\sum^{{\mathsf{K}}-D-1}_{i=1}{\mathbf{v}}^{[i]}\sum^{min({\mathsf{K}}-U-D,i)}_{s=max(1,i-U)}H^{{\mathsf{K}}+1,s}W_{i-s+1}^{[s]}$
@@
The interference dimension from communication signals is the same as that for communication users, i.e., ${\mathsf{K}}-D-1$. Therefore, we can get a left null vector ${\mathbf{v}}_{0}$ in $\begin{bmatrix}{\mathbf{v}}^{[D+2]}\ldots{\mathbf{v}}^{[{\mathsf{K}}]}\end{bmatrix}$ to remove the communication interference within ${\mathsf{K}}-D+U$ time slots, thus
${\mathbf{v}}_{0}{\mathbf{y}}^{[{\mathsf{K}}+1]}$ $=\negmedspace\sum_{i=D+2}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0}{\mathbf{x}}^{[i]}\negmedspace+\negmedspace{\mathbf{v}}_{0}{\mathbf{z}}^{[{\mathsf{K}}+1]}$
$=\sum_{i=D+2}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[i]}+z^{\prime[{\mathsf{K}}+1]}.$
Let $x_{1}^{[0]}=\begin{bmatrix}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[D+2]}&\cdots&{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[K-U]}\end{bmatrix}^{H}$, in other $({\mathsf{K}}-D-U)$ transmission, we can use the same way to get $x_{2}^{[0]},x_{3}^{[0]},\cdots,x_{{\mathsf{K}}-D-U}^{[0]}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\cdots,x_{{\mathsf{K}}-D-U}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\cdots,x_{{\mathsf{K}}-D-U}^{[0]}]^{H}=I$. So one effective observation of the channel can be obtained, achieving the sDoF equal to $\frac{1}{{\mathsf{K}}-D+U}$.
Appendix C Proof of Theorem 5-b: Topology 2 with Cooperative Transmitters
This is the general description of the proposed TIM scheme for Theorem 5-b. We then consider the bistatic ISAC systems with heterogeneous connectivity and cooperative transmitters. For the Topology 2 based on the $({\mathsf{K}},d)$-regular network, we propose an ISAC scheme based on TIM achieving the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{1}{d+1},\frac{2{\mathsf{K}}}{d+1}\right)$.

### v5

@@ VI Simulation Results
TABLE I: Simulation Parameters
Parameter Value
SNR range (dB) [5, 35]535[-5,\ 35][ - 5 , 35 ] (step 5 dB)
Number of transmitted symbols 1000
Transmit power (dBm) $30$ dBm
Path loss exponent 3.5
Number of Monte Carlo simulations $100$
Carrier frequency $2.4$ GHz
Sampling frequency $20$ MHz
Symbol rate $1$ MHz
LDPC code rate $0.5$
Samples per symbol $20$
Modulation order $4$ (DQPSK)
Rolloff factor $0.25$
Filter span (symbols) $6$
Comm. user distances Uniformly $[50,100]$ m
Sensing target distances Uniformly $[10,50]$ m
We simulate proposed schemes for the bistatic ISAC systems with heterogeneous coherent times and the interference channel in Example 1, and with heterogeneous connectivity and the interference channel in Example 2. The simulation flowchart is illustrated in Fig. 7 and the simulation parameters are provided in Table I. All the source codes can be found in [41].
[…]
Generation of the dedicated sensing signal. We set the power of the dedicated sensing signal $P_{s}=P/2$ as half of the total power. As shown in Footnote 8, since the statistic CSI is known at the ${\mathsf{m}}^{\prime}$ single antenna transmitters connected to the sensor, the transmitted deterministic sensing signal can be $\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}={\mathbf{W}}{\mathbf{S}}_{D}$ with a deterministic training signal ${\mathbf{S}}_{D}\in\mathbb{C}^{{\mathsf{m}}^{\prime}\times N}$ satisfying ${\mathbf{S}}_{D}{\mathbf{S}}_{D}^{H}=\mathbf{I}_{{\mathsf{m}}^{\prime}}$ and a precoding matrix ${\mathbf{W}}\in\mathbb{C}^{{\mathsf{m}}^{\prime}\times{\mathsf{m}}^{\prime}}$ to minimize detection/estimation errors. $\mathbf{W}$ is obtained by solving the following optimization problem based on the Linear Minimum Mean Square Error (LMMSE) criterion [42], where the goal is to minimize the mean squared error (MSE) subject to a transmit power constraint:
$\min_{\mathbf{W}\in\mathcal{A}}J_{\text{LMMSE}}=\text{tr}\left[\left(\mathbf{R}_{H}^{-1}+\frac{1}{\sigma^{2}}\mathbf{W}\mathbf{W}^{H}\right)^{-1}\right],$ (15)
where $\mathbf{R}_{H}$ denotes the statistical channel correlation matrix. By performing eigendecomposition of $\mathbf{R}_{H}=\mathbf{Q}\mathbf{\Lambda}\mathbf{Q}^{H}$ and applying the water-filling principle, the optimal solution can be expressed as:
$\mathbf{W}=\sqrt{\frac{\sigma^{2}}{N}}\,\mathbf{Q}\left[\left(\mu_{0}\mathbf{I}_{{\mathsf{m}}^{\prime}}-\mathbf{\Lambda}^{-1}\right)^{+}\right]^{\frac{1}{2}},$ (16)
where $\mathbf{\Lambda}^{-1}=\text{diag}(\lambda_{1}^{-1},\lambda_{2}^{-1},\ldots,\lambda_{m^{\prime}}^{-1})$, $(\cdot)^{+}$ denotes the positive part function $(\mu_{0}-\lambda_{i}^{-1})^{+}=\max(\mu_{0}-\lambda_{i}^{-1},0)$, and the water level $\mu_{0}$ is chosen to satisfy the power constraint:
$\frac{\sigma^{2}}{N}\sum_{i=1}^{{\mathsf{m}}^{\prime}}(\mu_{0}-\lambda_{i}^{-1})^{+}=P_{s}.$ (17)
Generation of the dedicated communication signal. Once the sensing signal is precoded, it is superimposed with the modulated communication signal in the time domain. We set the power of the dedicated communication signal $P_{c}=P/2$, also as half of the total power. The resulting baseband signal is oversampled at a rate of 20 samples per symbol. The system employs Low-Density Parity-Check (LDPC) code channel coding followed by DQPSK modulation to encode the communication bits. The LDPC code rate is $0.5$, and the symbol rate is set to $1$ MHz, corresponding to a sampling frequency of $20$ MHz. Pulse shaping is applied using a raised cosine filter with a roll-off factor of $0.25$ and a span of $6$ symbols. The carrier frequency is fixed at $2.4$ GHz.
Benchmark schemes. For the sake of comparison, we consider two benchmark schemes that differ in how the sensor handles the communication signal, where the TIN scheme treats the communication signal as noise without decoding, and the SIC scheme fully decodes the communication messages and then removes them to improve sensing performance. At the transmitter side, both benchmark schemes directly send the sum of the dedicated sensing and communication signals, while the proposed schemes additionally precode the dedicated sensing signal and communication signals based on BIA or TIM.
For both of benchmark schemes, the communication receiver directly removes the interference of the dedicated sensing signal and then decodes the message. Let us then introduce the main steps of the benchmark schemes at the sensor side. The TIN benchmark scheme consists of the following steps:
  1. 1.
    Matched filtering and sampling: The received ISAC waveform undergoes downconversion, matched filtering (using a square-root raised cosine filter), and symbol-rate sampling to extract the complex baseband symbols. In this method, interference from communication symbols is not mitigated but instead treated as noise.
  2. 2.
    Channel estimation: Using estimate method to estimate the channel.
The SIC benchmark scheme consists of the following steps:
  1. 1.
    Matched filtering and sampling: The received ISAC signal is downconverted and passed through a matched filter (e.g., square-root raised cosine). It is then sampled at the symbol rate to extract complex-valued baseband symbols.
  2. 2.
    Differential phase detection: For each extracted symbol $y_{n}$, the phase difference with its predecessor is computed as
    $\Delta\phi_{n}=\angle(y_{n})-\angle(y_{n-1}),$
    followed by normalization into the range $(-\pi,\pi]$ to account for phase wrapping.
  3. 3.
    Nearest constellation matching: The normalized phase difference $\Delta\phi_{n}$ is compared against the ideal DQPSK phase shifts $\{0,\pi/2,\pi,3\pi/2\}$. The detected symbol index is determined by selecting the closest phase:
    $\hat{a}_{n}={\hbox{arg}}\min_{k\in\{0,1,2,3\}}\left|\Delta\phi_{n}-\theta_{k}\right|\mod 2\pi,$
    where $\theta_{k}=\frac{2\pi k}{M}$, and $M=4$ for DQPSK.
  4. 4.
    Symbol remapping: The detected indices $\hat{a}_{n}\in\{0,1,2,3\}$ are mapped to complex DQPSK constellation points .
  5. 5.
    ISAC signal reconstruction and channel estimation: The recovered communication symbols and sensing signals are used to reconstruct the ISAC signal, enabling subsequent radar channel estimation.
Figure 8: Simulation result of BIA.
Figure 9: Simulation result of TIM.
@@ VI-A Heterogeneous Coherence Times
For the bistatic ISAC systems with heterogeneous coherent times and the interference channel in Example 1, as shown in Fig. 8, the proposed method demonstrates superior robustness to interference, especially in the high-SNR regime. It significantly reduces the CEE as SNR increases. In contrast, the TIN benchmark scheme treats the communication signal as noise, which dominates sensing performance at high power levels, resulting in persistently high estimation error even as noise decreases.
Quantitatively, our method achieves an estimation accuracy gain of $[7.83,47.81]$ dB over the TIN benchmark scheme. Compared with the SIC benchmark scheme which performs noncoherent detection followed by signal subtraction—the proposed scheme also shows consistent improvement. While the SIC benchmark scheme suffers from poor detection at low SNR and improves at higher SNR, our method still outperforms it by $[7.17,7.85]$ dB in the high-SNR regime.
The proposed scheme exhibits an overall complexity of $O({\mathsf{K}}^{3})$, primarily due to the LS-based channel estimation, with an additional differencing step of only $O({\mathsf{K}})$ complexity. The SIC benchmark scheme also incurs $O({\mathsf{K}}^{3})$ complexity, comprising $O({\mathsf{K}}^{2})$ for noncoherent detection and $O({\mathsf{K}}^{3})$ for channel estimation, along with added overhead from signal reconstruction. The TIN benchmark scheme skips symbol decoding and directly performs channel estimation, resulting in a simpler structure but the same asymptotic complexity of $O({\mathsf{K}}^{3})$. Thus, while all methods share the same complexity order, their practical implementation overheads differ.
It is important to highlight that all three schemes yield comparable communication performance (e.g., in terms of rate or BER). In other words, with the same communication performance, our proposed method significantly improves the sensing performance.
@@ VI-B Heterogeneous Connectivity
For the bistatic ISAC systems with heterogeneous connectivity and the interference channel in Example 2, as shown in Fig. 9, compared to the TIN benchmark scheme, our proposed approach achieves a significant improvement in channel estimation accuracy, with a gain ranging from $[7.49,48.19]$ dB. In addition, compared to the SIC benchmark scheme, our approach still exhibits a performance advantage of $[7.65,7.99]$ dB under high SNR conditions. As in the scenario of heterogeneous coherence times, the SIC benchmark scheme performs poorly in low SNR regimes due to residual interference and limited noise suppression capability, and as the SNR increases, it gradually improves and becomes more effective in mitigating communication-induced interference, thus reducing the channel estimation error.
The proposed scheme incurs a total complexity of $O\left(({\mathsf{K}}+U)\cdot L^{2}+L^{3}\right)$ with $L={\mathsf{K}}-D+U+1$, where the first term arises from precoding construction and the second from sensing signal design and LS-based channel estimation. The SIC benchmark scheme involves noncoherent detection and LS estimation, with overall complexity $O(L^{3})$, while the TIN benchmark scheme bypasses symbol detection and relies solely on LS estimation, also at $O(L^{3})$. Although all methods share the same asymptotic complexity order with respect to $L$, the proposed scheme requires additional steps for structured signal design, whereas the TIN benchmark scheme offers the lowest implementation overhead.
Remark 4 (Discussion on the theoretic assumptions).
The theoretical results in Section III are established under the following common assumptions: (1) the transmitters have perfect CSI or knowledge of the network topology of the communication channel, and the communication receivers have perfect CSIR; (2) the communication and sensing channels have heterogeneity, either in coherence time or in network topology; (3) the tradeoff between sDof and cDof is considered.
However, in our simulations, we discard assumption (3) by considering the practical sensing and communication metrics. It can be seen from the simulation results that, our interference elimination strategies based on BIA and TIM can also improve the ISAC performance on practical metrics (e.g., estimation error, communication rate/error).
Fundamentally, all our schemes exploit the heterogeneity inherent in ISAC systems—such as differences in coherence time and network connectivity structures—to enable effective interference elimination and performance enhancement. In the ongoing work, we are extending this interference elimination strategy to more general ISAC systems, under more relaxed assumptions.
@@ VII Conclusion
This paper explored the interference management issue in bistatic ISAC systems, specifically how to effectively cancel the interference from the communication signals to the sensor, by applying the BIA and TIM strategies. The key is to leverage the heterogeneity (on the coherence times or on the connectivity) between the sensor and receivers. By using the communication and sensing degrees of freedom as the metric, the proposed schemes provide improved tradeoff points, compared to the time-sharing between the communication-only and sensing-only points. Simulations on the proposed schemes and on the existing TIN, SIC schemes are provided, showing that the proposed schemes reduce the sensing CEE, while maintaining the same communication performance.
@@ Appendix A Proof of Theorem 3: MU-MIMO
Appendix B Proof of Remark 1
B-A Proof of (7)
This is the general description of the proposed TIM scheme for the $({\mathsf{K}},U,D)$ neighboring antidotes network with an adding sensor. For this framework, we propose an ISAC scheme based on TIM achieving the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{U+1}{{\mathsf{K}}-D+U},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}\right)$ in (7). Each receiver is disconnected from the previous $U$ and next $D\geq U$ transmitters, and connected to other ${\mathsf{K}}-U-D>0$ transmitters. Remind that the sensor is a subset of one communication that the network topology could be pressed as
@@ B-A Proof of (7)
for each $i,k\in\{1,\ldots,{\mathsf{K}}\}$, To simplify notation without loss of generality, we set $i=1$, thus we can get ${\mathcal{R}}_{s}=\{2+D\,\ldots,{\mathsf{K}}-U\}$. Let $t_{0}=({\mathsf{K}}-D+U)\left\lceil\frac{({\mathsf{K}}-D+U)}{U+1}\right\rceil$, during these $({\mathsf{K}}-D+U)\left\lceil\frac{({\mathsf{K}}-D+U)}{U+1}\right\rceil$ time slots, the sensor will obtain $(U+1)\left\lceil\frac{({\mathsf{K}}-D+U)}{U+1}\right\rceil=N$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}\in{\mathbb{C}}^{({\mathsf{K}}-D-U-1)\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}]^{H}=I$. In each period of $({\mathsf{K}}-D+U)$ time slots, we let each receiver decode $U+1$ communication messages, and let the sensor obtain $U+1$ effective observations. In the next, we will illustrate our proposed scheme for the first $K-D+U$ time slots, where each receiver $k\in\{1,2,...,{\mathsf{K}}\}$ should recover $U+1$ messages $W_{1}^{[k]},\ldots,W_{U+1}^{[k]}\in{\mathbb{C}}$.
@@
for each $i\in{\mathcal{R}}_{s}$, where ${\mathbf{x}}_{s}^{[j]},j\in\mathcal{R}_{s}$ are the sensing signals to be determined later. For the transmitters not connected to the sensor, design based on the cyclic coding,
@@
where $\max(a,b)$ is the larger value (i.e., $a$ if $a\geq b$, otherwise $b$) and $\min(a,b)$ is the smaller value (i.e., $a$ if $a\leq b$, otherwise $b$). Therefore, the dimension of the interference signals is aligned to ${\mathsf{K}}-D-1$, by the linear independence of ${\mathsf{K}}-D-1$ vectors ${\mathbf{v}}^{[<k+D+1>_{{\mathsf{K}}+1}]}\ {\mathbf{v}}^{[<k+D+2>_{{\mathsf{K}}+1}]}\ \ldots{\mathbf{v}}^{[<k+{\mathsf{K}}>_{{\mathsf{K}}+1}]}$, receiver $k$ can recover $U+1$ messages from the first ${\mathsf{K}}-D+U$ time slots. Thus, the communication degree of freedom is $\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}$.
@@
$=\sum^{{\mathsf{K}}-D-1}_{i=1}{\mathbf{v}}^{[i]}\sum^{min({\mathsf{K}}-U-D,i)}_{s=max(1,i-U)}H^{[{\mathsf{K}}+1,s]}W_{i-s+1}^{[s]}$
@@
The interference dimension from communication signals is the same as that for communication users, i.e., ${\mathsf{K}}-D-1$. Therefore, we can get $U+1$ left null vectors ${\mathbf{v}}_{0,1},{\mathbf{v}}_{0,2},\dots,{\mathbf{v}}_{0,U+1}$ in $\begin{bmatrix}{\mathbf{v}}^{[D+2]}\ldots{\mathbf{v}}^{[{\mathsf{K}}]}\end{bmatrix}$ to remove the communication interference within ${\mathsf{K}}-D+U$ time slots, thus
${\mathbf{v}}_{0,u}{\mathbf{y}}^{[{\mathsf{K}}+1]}$ $=\negmedspace\sum_{i=D+2}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0,u}{\mathbf{x}}^{[i]}\negmedspace+\negmedspace{\mathbf{v}}_{0,u}{\mathbf{z}}^{[{\mathsf{K}}+1]}$
$=\sum_{i=D+2}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[i]}+z_{u}^{\prime[{\mathsf{K}}+1]}.$
Let $x_{u}^{[0]}=\begin{bmatrix}{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[D+2]}&\cdots&{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[{\mathsf{K}}-U]}\end{bmatrix}^{H}$, in other $({\mathsf{K}}-D-U)$ transmission, we can use the same way to get $x_{U+2}^{[0]},x_{U+3}^{[0]},\cdots,x_{{\mathsf{K}}-D-U}^{[0]}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\cdots,x_{{\mathsf{K}}-D-U}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\cdots,x_{{\mathsf{K}}-D-U}^{[0]}]^{H}=I$. In other words, we solve ${\mathbf{x}}_{s}^{[D+2]}$ with ${\mathsf{K}}-D+U$ parameters satisfying ${\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[D+2]}=x_{u,1}$, where $x_{u,1}$ is the first element of ${\mathbf{x}}_{u}^{[0]}$, and similarly for ${\mathbf{x}}_{s}^{[D+3]},\dots,{\mathbf{x}}_{s}^{[{\mathsf{K}}-U]},{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}$. So one effective observation of the channel can be obtained, achieving the sDoF equal to $\frac{U+1}{{\mathsf{K}}-D+U}$.
B-B Proof of (8)
This is the general description of the proposed TIM scheme for the $({\mathsf{K}},d)$ regular network with an adding sensor. In this framework, we propose an ISAC scheme based on TIM achieving the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{1}{d+1},\frac{2{\mathsf{K}}}{d+1}\right)$.